OpenStudy (anonymous):

PLEASE HELP. !!! Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (merchandize):

m=-1 x=-4 y=2 y=mx+c 2=-1*(-4)+c c=-2 so y=-x-2